The question:
Suppose that the length of a rectangle is one and one-third times as long as its width. The area of the rectangle is 48 square centimeters. Find the length and width of the rectangle.
I need help devising/solving the equation:
a=lw
48=(1 and 1/3)(L)(W)

Answer by venugopalramana(3286)

Suppose that the length of a rectangle is one and one-third times as long as its width. The area of the rectangle is 48 square centimeters. Find the length and width of the rectangle.
I need help devising/solving the equation:
a=lw...VERY GOOD
48=(1 and 1/3)(L)(W)...ALMOST CORRECT..THE BEST WAY IS
LET WIDTH =W
LENGTH=one and one-third times as long as its width.=(1 and 1/3)(W)..=(4/3)(W).
=4W/3
SO AREA =L*W=(4W/3)*W=48
W*W=48*3/4=36
W=6 CM
L=4W/3=4*6/3=8 CM